Q. $A$ and $B$ are two independent events such that $P\left( A \cup B^{'} \right) = 0.8$ and $P(A) = 0.3.$ Then, $P(B)$ is

A) $\frac{2}{7}$
B) $\frac{2}{3}$
C) $\frac{3}{8}$
D) $\frac{1}{8}$
Explanation

(a)
$P\left( A \cup B^{'} \right) = P(A) + P\left( B^{'} \right) - P(A)P(B^{'})$
$\therefore 0.8 = 0.3 + P\left( B^{'} \right) - 0.3P(B')$
$\Rightarrow 0.5 = P\left( B^{'} \right)(0.7)$
$\Rightarrow P\left( B^{'} \right) = \frac{5}{7}$
$\therefore P(B) = 1 - \frac{5}{7} = \frac{2}{7}$

Dr. Abubacker
Reviewed by Dr. Abubacker

Physics Educator & NEET/JEE Specialist

Ph.D. in Physics with 19+ years of teaching experience.